OpenStudy (jkbo):

please help A salesperson earns a salary of $800 per month plus 2% of the sales. Which inequality represents the total sales if the salesperson is to have a monthly income of at least $1600?

OpenStudy (jdoe0001):

it helps ifyou provide us with the inequalities options :/

OpenStudy (jkbo):

A. x > $40,000 B. x < $36,000 C. x > $32,000 D. x < $44,000

OpenStudy (anonymous):

We want $800 + 2% of the sales to be greater than $1600. If we denote the sales with x, we have: \[800+\frac{ 2 }{ 100 }x > 1600\] Subtracting 800 from both sides: \[\frac{ 2 }{ 100 }x > 800\] Dividing both sides by 2: \[\frac{ 1 }{ 100 }x > 400\] Multiplying both sides by 100: \[x > 400 *100 \] Making the 400*100 multiplication \[x > 40000\] So the correct answer would be A. (In my opinion income of at least $1600 denotes greater or equal operator, but if those are the possible answers, that's ok.)